OpenStudy (anonymous):

Find the equation, in standard form, of the line passing through the points (2,-3) and (4,2).

OpenStudy (nincompoop):

let us do this step-by-step okay? I will give you hints and you the rest

OpenStudy (nincompoop):

first, solve for the slope

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the slope is 5/2, right?

OpenStudy (nincompoop):

\[slope=m=\frac{ \Delta y }{ \Delta x }=\frac{ rise }{ run }=\frac{ y_2 - y_1 }{ x_2 - x_1 }\]

OpenStudy (anonymous):

Ok, i did that. I did 2-(-3)/ 4-2 and i got 5/2

OpenStudy (nincompoop):

now that you have the slope, use the point-slope form equation \[y - y_1 = m(x-x_1)\]

OpenStudy (anonymous):

ooh

OpenStudy (nincompoop):

go head, plug in the numbers

OpenStudy (anonymous):

y-2=5/2(x-4)?

OpenStudy (nincompoop):

use distributive property to multiply the slope into x and -4

OpenStudy (anonymous):

y-2=5/2x-10

OpenStudy (nincompoop):

not 5/2x it is 5x/2 be very careful

OpenStudy (anonymous):

ah

OpenStudy (anonymous):

add 2 to both sides?

OpenStudy (nincompoop):

\[\frac{ 5 }{ 2 } \times \frac{ x }{ 1 } = \frac{ 5x }{ 2 }\]

OpenStudy (nincompoop):

then solve for y you can achieve this by isolating y and in your problem you need to add 2 on both sides of the equation (left and and right hand sides)

OpenStudy (anonymous):

and that's the answer in standard form? y=5x/2-8?

OpenStudy (nincompoop):

nope that is the slope form

OpenStudy (nincompoop):

standard form is \[Ax+By=C\]

OpenStudy (anonymous):

so would it be 5x+2y=8?

OpenStudy (nincompoop):

how so?

OpenStudy (anonymous):

because the 5 is next to x, the -8 is constant and 2 is what's left i guess? and i'm not sure if they can be negative

OpenStudy (nincompoop):

first, rearrange it in the format of Ax + By = C the coefficient of x is 5/2 the coefficient of y is 1 and C = 8 rearranging \[y = \frac{ 5 }{ 2 }x - 8\] add 8 to both sides subtract y to both sides then make sure that everything is in whole number

OpenStudy (nincompoop):

always make sure that A is a positive integer

OpenStudy (anonymous):

ok, so is it 5x-2y=16?

OpenStudy (nincompoop):

yes

OpenStudy (anonymous):

yay, thank you so much :)!